# 3D Shape Generation: A Survey

### v1

@@ Abstract
Recent advances in deep learning have significantly transformed the field of 3D shape generation, enabling the synthesis of complex, diverse, and semantically meaningful 3D objects. This survey provides a comprehensive overview of the current state of the art in 3D shape generation, organizing the discussion around three core components: shape representations, generative modeling approaches, and evaluation protocols. We begin by categorizing 3D representations into explicit, implicit, and hybrid setups, highlighting their structural properties, advantages, and limitations. Next, we review a wide range of generation methods, focusing on feedforward architectures. We further summarize commonly used datasets and evaluation metrics that assess fidelity, diversity, and realism of generated shapes. Finally, we identify open challenges and outline future research directions that could drive progress in controllable, efficient, and high-quality 3D shape generation. This survey aims to serve as a valuable reference for researchers and practitioners seeking a structured and in-depth understanding of this rapidly evolving field.
1 Introduction
Generative modeling of 3D shapes has gained significant attention due to its broad applicability across diverse domains, including medicine [GYS+18, FWB+23, DWB+25], robotics [NZZC20], bioinformatics [HSVW22, XYS+21, YZP+23], computer-aided design [HPG+22, LLW22, HHL+24b], cultural heritage [HS18, XSG+24], and more. To support such downstream tasks, a successful 3D generative model should be both faithful and probabilistic [ZDW21]. A faithful model generates geometrically plausible and visually realistic shapes, while also respecting partial observations (e.g., depth maps or occlusions) when conditioned. On the other hand, a probabilistic model must capture the inherent ambiguity and multimodality of 3D shape generation, allowing diverse outputs from incomplete or ambiguous inputs.
Given the data-intensive nature of deep learning, the construction and benchmarking of large-scale 3D shape datasets have become essential for training and evaluating generative models [CFG+15, ZSK+15].
However, learning to generate 3D shapes introduces unique challenges. Thus, the complexity and irregularity of 3D shapes and the different requirements of practical applications have resulted in the lack of a unified representation of 3D shapes [XLZ+20]. Instead, various 3D representations are employed, each with its own structural properties, advantages, and limitations. These factors directly impact model design, data processing, and evaluation strategies.
We organize the discussion around three core components of the 3D shape generation pipeline: representations, generation methods, and evaluation protocols. Figure 1 presents an overview of this survey. We first discuss the scope and related works in Section 2. Section 3 introduces major categories of 3D shape representations: explicit, implicit, and hybrid. Section 4 explores the most relevant 3D shape generation methods, with a focus on the feedforward generative models. Sections 5 and 6 present widely used datasets and metrics for assessing the fidelity and diversity of generated shapes, respectively. Finally, in Section 7, we discuss the open challenges and future directions.
@@ 1 Introduction
3DGAN [WZX+16] GAN Latent 1D Voxel ShapeNet, ModelNet, IKEA Average precision
@@
SDM-Net [GYW+19] VAE Latent 1D Mesh ShapeNet, ModelNet JSD, MMD, Cov, CD, EMD
TM-Net [GWY+21] VAE Latent 1D Mesh ShapeNet, ModelNet Image similarity measures
@@
Michelangelo [ZLC+23] VAE Latent 1D Mesh ShapeNet IoU, Other proposed by the authors
PointFlow [YHH+19] Normalizing Flow Latent 1D Point Cloud ShapeNet, ModelNet JSD, MMD, Cov, 1-NNA, CD, EMD
@@
SoftFlow [KLK+20] Normalizing Flow Latent 1D Point Cloud ShapeNet 1-NNA, CD, EMD
@@
PointGrow [SWL+20] Autoregressive Latent 3D Point Cloud ShapeNet, ModelNet, PASCAL3D+ PointNet Distance (PND - Derived from FID)
@@
PointARU [MWZ+25] Autoregressive Latent 3D Point Cloud ShapeNet 1-NNA, CD, EMD
@@
LION [ZVW+22] Diffusion Latent 3D Point Cloud ShapeNet 1-NNA, Cov, MMD, CD, EMD
@@
SLIDE [LWA+23] Diffusion Latent 3D Point Cloud, Mesh ShapeNet 1-NNA, MMD, Cov, CD, EMD
DiffTF [CHW+23] Diffusion Latent 3D Triplane ShapeNet, OmniObject FID, KID, Cov, MMD, CD
Sin3DM [WLVZ23] Diffusion Latent Diffusion Triplane Single 3D Tectured Shape G-Qual, SSFID, G-Div, IoU
XCube [RHZ+24] Diffusion Latent Diffusion Voxel ShapeNet, Objaverse 1-NNA, CD, EMD
@@
GPLD3D [DZG+24] Diffusion Latent Diffusion Mesh ShapeNet FPD, KPD
Direct3D [WLZ+24] Diffusion Latent Diffusion Mesh Objaverse, GSO CD, Volume IoU, F-Score
@@
Figure 1: Overview of this survey, covering 3D representations, generation methods, datasets, and evaluation metrics. We first introduce the 3D representations that serve as the foundation for 3D shape generation. Next, we provide a comprehensive overview of the rapidly expanding body of literature on generation methods, focusing specifically on feedforward paradigms. Finally, we briefly describe the most widely used datasets for 3D generation and the metrics employed to evaluate the quality of the generated shapes. Source [LZK+24].
@@ 2 Scope of this Survey
This survey aims to systematically summarize and categorize methods for 3D shape generation, together with the relevant datasets and metrics. The surveyed papers are primarily published in top-tier computer vision and computer graphics conferences and journals, as well as selected preprints published on arXiv in 2025. While it is challenging to exhaustively cover all methods related to 3D shape generation, we strive to include the most significant branches of the field. Rather than providing exhaustive explanations for each branch, we highlight representative works (as shown in Figure 2) to illustrate their main contributions and limitations. Further details can be found in the related work of these cited papers.
@@ 2.1 Related Survey
Our primary focus is on exploring techniques that generate 3D shapes. Therefore, this review does not encompass research on generation methods in the 2D domain. Several recent surveys have addressed specific aspects of 3D generation, including Deep Learning for 3D point cloud [GWH+21], 3D-aware image synthesis [XX23], 3D generative models [SPX+23], Text-to-3D [LZC+24], 3D generation methods based on scene representation [LZK+24], implicit representations-based 3D shape generation [SWG24], and diffusion models for 3D generation [WPL+25, Kum25]. In contrast, this survey provides a comprehensive review of 3D shape generation methods and their related datasets and metrics (as shown in Table 1).
@@
3D shape representations are central to 3D computer vision and graphics [PJL+21]. Shape representations can generally be categorized as either explicit or implicit. Explicit shape representations and learning algorithms depending on such representations directly parameterize the surface of the geometry, either as a point cloud [FSG17a, MYTB21, QSMG17, YHH+19, WWG+23, MWZ+25], meshes [WZL+18, JHK+19, GC20], or surface patches [GFK+18, WSS+19].
While explicit representations are often computationally efficient and require few parameters, they tend to suffer from several limitations: discretization artifacts, difficulties in representing watertight surfaces, and constraints imposed by fixed or pre-defined topologies (meshes). In contrast, implicit representations define shapes as level sets of continuous functions over 3D space. These functions can be represented using discretized voxel grids [DRB+18, WZX+16] or, more recently, parameterized using neural networks, commonly referred to as neural implicit functions [YCWW22, BRS+23].
@@ 3.1.1 Point Clouds
According to [PFS+19], a point cloud is a lightweight 3D representation that closely matches the raw data that many sensors (e.g., LiDARs, depth cameras) provide, and hence is a natural fit for applying 3D learning.
@@
A primary limitation of point clouds is that they do not describe the topology and are not suitable for producing watertight surfaces [PFS+19]. Another issue with point clouds as a representation is that they are unordered, which means that any permutation of a point set still describes the same shape. This complicates comparisons between two point sets, typically needed to define a loss function [ADMG18].
@@ 3.1.2 Meshes
According to [BKP+10, BBL+17, WZL+18], a 3D mesh is a collection of vertices, faces, and edges that defines the shape of a 3D object; it can be represented by a graph structure (simplicial complex) $\mathcal{M}=(\mathcal{V},\mathcal{F},\mathcal{E})$, where
@@
=$\mathcal{F}=\{f_{1},\dots,f_{F}\},\quad f_{i}\in\mathcal{V}\times\mathcal{V}\times\mathcal{V}$
is a set of triangular faces connecting them. However, it is more efficient to represent the connectivity of a triangle mesh in terms of the edges of the respective graph [BKP+10],
=$\mathcal{E}=\{e_{1},\dots,e_{E}\},\quad e_{i}\in\mathcal{V}\times\mathcal{V}.$
In this sense, a mesh-based representation stores the surface information cheaply as a list of vertices and faces that respectively define the geometric and topological information. Early approaches to mesh generation focused on predicting the parameters of category-specific mesh models [ZKB18, KPD19]. Although these methods produce manifold meshes, they are limited to object categories for which parameterized templates are available. Building upon this idea, works such as [GFK+18, WZL+18] extended mesh generation to a wider variety of object categories by incorporating topological priors, enabling more flexible and expressive 3D shape modeling beyond category-specific templates.
Compared to voxels, meshes are more compact and better suited to representing finer surface details. Compared to points, meshes are more controllable and exhibit better visual quality [GYW+19].
@@ 3.1.3 Voxel Grids
The direct extension of pixels in 2D images to 3D is the voxel representation, which has a regular grid structure well-suited for convolutional neural networks (CNNs) [GFRG16, WZX+16, GYW+19].
@@
Moreover, as highlighted by [DRB+18], small voxel sizes capture local detail but lack spatial context; large voxel sizes provide large spatial context but lack local detail.
@@ 3.2.1 Neural Radiance Fields (NeRF)
[MST+22] introduce NeRF, a method for representing a 3D scene as an implicit function defined as
@@
where $(\mathbf{x}$ is a 3D spatial coordinate, $\mathbf{d}$ is a 3D viewing direction, $\mathbf{c}$ is an RGB color, and $\sigma$ is a non-negative density value.
NeRFs model the appearance and geometry of 3D shapes by using volume rendering (ray matching) algorithms and positional encoding [MST+22, MRP+23]. This technique can efficiently reconstruct geometry given only a few multiview posed images. To achieve higher geometry quality, NeuS [WLL+21] and VolSDF [YGKL21] extend radiance fields by using SDF instead of density as the geometry [SWG24].
@@ 3.2.2 Signed Distance Field (SDF)
As a function, SDF is more flexible than the common explicit representations, such as point clouds or meshes, and inherently allows topology manipulations such as constructive solid geometry (CSG) operations. Moreover, SDF allows the use of a technique known as “sphere tracing” [Har96], which can accelerate the rendering of path tracing [SWG24].
@@ 3.2.3 Neural Implicit Functions (NIFs)
Neural implicit functions (NIFs) have recently gained significant attention as a powerful and flexible framework for representing 3D scenes using neural networks. Unlike traditional explicit representations such as meshes, NIFs are not constrained by fixed spatial resolution, allowing them to capture fine geometric details at arbitrary scales [BRS+23].
@@ 3.3.1 Point-Voxel
To address these limitations, [ILK21] proposes the use of point-voxel representations for 3D shape generation. This hybrid representation had been previously explored in methods such as PointCNN [LBS+18], PV-RCNN [SGJ+20], and PVD [ZDW21]. The core idea behind point-voxel is to voxelize the point clouds to enable efficient 3D convolutions while preserving point-level detail for finer geometric modeling.
@@ 3.3.2 Tri-Plane
In [CLC+22], the Tri-Plane representation was introduced as a hybrid explicit-implicit 3D representation that balances computational efficiency with representational expressiveness.
In this formulation, features are explicitly arranged on three orthogonal, axis-aligned 2D planes–namely the $xy$, $yz$, and $xz$ planes, each with a resolution $N\times N\times C$, where $N$ denotes spatial resolution and $C$ the number of feature channels. As illustrated in Fig. 3, this structure contrasts with traditional 3D representations by encoding volumetric information through multiple planar projections.
Figure 3: Comparison of neural implicit, voxel-based, and tri-plane representations. (a) Neural implicit representations use fully connected layers (FC) with positional encoding (PE), but are slow to query. (b) Explicit voxel grids and hybrid variants using local decoders are faster but scale poorly with resolution. (c) The tri-plane representation enables efficient queries and better scaling with resolution. Source [CLC+22].
@@
In contrast to the original formulation, [SCP+23] confirmed that simple addition (as shown in Equation 2) is an effective and efficient aggregation strategy for tri-plane features, avoiding the complexity of more elaborate fusion mechanisms.
The primary advantage of tri-planes representation lies in its efficiency–by keeping the decoder small and shifting the bulk of the expressive power into the explicit features. As demonstrated by [CLC+22], this approach reduces the computational cost of neural rendering compared to fully implicit MLP architectures without losing expressiveness.
@@
According to [KYNS23], the first 3D generative models are based on GANs, learning a distribution of latents that can be decoded into various 3D representations such as point clouds [ADMG18, VFM18, SPK19] and implicit representations [KFW20, HASB20, CZ19, ILK21, ZLWT22]. Building upon this foundation, the rapid development of generative models introduces new generation methods such as flow-based models [YHH+19, KBV20], which learn invertible mappings between complex and straightforward point distributions, enabling exact likelihood estimation and diverse point cloud synthesis; variational autoencoders (VAEs) [GFK+18], which encode 3D shapes into a probabilistic latent space and reconstruct them via learned decoders, allowing interpolation and latent sampling; autoregressive models [SWL+20, MGY+19], which generate 3D geometry sequentially–either point-by-point or part-by-part–capturing structural dependencies; and, more recently, diffusion-based approaches [ZVW+22, NJD+22], which progressively refine noisy inputs into clean 3D shapes through iterative denoising steps. Table 1 shows well-known examples of 3D generation methods using generative models. Table 1 also reports the metrics and datasets used to evaluate each generation method.
In this section, we explore the feedforward generation paradigm, proposed by [LZK+24], which enables the direct synthesis of 3D representations through generative models.

### v2

@@ Abstract
Recent advances in deep learning have significantly transformed the field of 3D shape generation, enabling the synthesis of complex, diverse, and semantically meaningful 3D objects. This survey provides a comprehensive overview of the current state-of-the-art in 3D shape generation, organizing the discussion around three core components: shape representations, generative modeling approaches, and evaluation protocols. We begin by categorizing 3D representations into explicit, implicit, and hybrid setups, highlighting their structural properties, advantages, and limitations. Next, we review a wide range of generation methods, focusing on feedforward architectures. We further summarize commonly used datasets and evaluation metrics that assess fidelity, diversity, and realism of generated shapes. Finally, we identify open challenges and outline future research directions that could drive progress in controllable, efficient, and high-quality 3D shape generation. This survey aims to serve as a valuable reference for researchers and practitioners seeking a structured and in-depth understanding of this rapidly evolving field. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010178.10010224</concept_id> <concept_desc>Computing methodologies Computer vision</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
[500]Computing methodologies Computer vision
[…]
@@
3DGAN [WZX16] GAN Latent 1D Voxel ShapeNet, ModelNet, IKEA Average precision
@@
SDM-Net [GYW19] VAE Latent 1D Mesh ShapeNet, ModelNet JSD, MMD, Cov, CD, EMD
TM-Net [GWY21] VAE Latent 1D Mesh ShapeNet, ModelNet Image similarity measures
@@
Michelangelo [ZLC23] VAE Latent 1D Mesh ShapeNet IoU, Other proposed by the authors
TAR3D [ZLL25] VAE Latent 3D Triplane ShapeNet, Objaverse, GSO PSNR, LPIPS, SSIM, CLIP
COD-VAE [CYJK25] VAE Latent 1D Triplane ShapeNet, Objaverse IoU, CD, F-Score, FPD, MMD, Cov, 1-NNA
PointFlow [YHH19] Normalizing Flow Latent 1D Point Cloud ShapeNet, ModelNet JSD, MMD, Cov, 1-NNA, CD, EMD
@@
SoftFlow [KLK20] Normalizing Flow Latent 1D Point Cloud ShapeNet 1-NNA, CD, EMD
@@
PointGrow [SWL20] Autoregressive Latent 3D Point Cloud ShapeNet, ModelNet, PASCAL3D+ PointNet Distance (PND - Derived from FID)
@@
PointARU [MWZ25] Autoregressive Latent 3D Point Cloud ShapeNet 1-NNA, CD, EMD
ShapeGPT [YCZ25] Autoregressive Latent 1D Mesh ShapeNet IoU, CD, F-Score, P-FID, ULIP, CLIP
@@
LION [ZVW22] Diffusion Latent 3D Point Cloud ShapeNet 1-NNA, Cov, MMD, CD, EMD
@@
SLIDE [LWA23] Diffusion Latent 3D Point Cloud, Mesh ShapeNet 1-NNA, MMD, Cov, CD, EMD
DiffTF [CHW23] Diffusion Latent 3D Triplane ShapeNet, OmniObject FID, KID, Cov, MMD, CD
Sin3DM [WLVZ23] Diffusion Latent 3D Triplane Single 3D Textured Shape G-Qual, SSFID, G-Div, IoU
XCube [RHZ24] Diffusion Latent Diffusion Voxel ShapeNet, Objaverse 1-NNA, CD, EMD
@@
GPLD3D [DZG24] Diffusion Latent Diffusion Mesh ShapeNet FPD, KPD
Direct3D [WLZ24] Diffusion Latent Diffusion Mesh Objaverse, GSO CD, Volume IoU, F-Score
@@
ShapeShifter [MYF25] Diffusion Latent 3D Voxel, Points Single 3D Textured Shape SSFID, IoU
CoPart [DDC25] Diffusion Latent 3D, 2D Mesh PartVerse CLIP, ULIP
Hunyuan3D 2.1 [Hea25] Diffusion Latent Diffusion Mesh ShapeNet, ModelNet40, Thingi10K, Objaverse ULIP, Uni3D
VolGen [TWG25] Diffusion Latent 3D Mesh Objaverse CD, IoU, ULIP, Uni3D
Hi3DGen [YWL25] Diffusion Latent Diffusion Mesh, Normal map DetailVerse Image-to-normal estimation metrics
1 Introduction
Generative modeling of 3D shapes has gained significant attention due to its broad applicability across diverse domains, including medicine [GYS18, FWB23, DWB25], robotics [NZZC20], bioinformatics [HSVW22, XYS21, YZP23], computer-aided design [HPG22, LLW22, HHL24b, LJK25], cultural heritage [HS18, XSG24], and more. To support such downstream tasks, a successful 3D generative model should be both faithful and probabilistic [ZDW21]. A faithful model generates geometrically plausible and visually realistic shapes, while also respecting partial observations (e.g., depth maps or occlusions) when conditioned. On the other hand, a probabilistic model must capture the inherent ambiguity and multimodality of 3D shape generation, allowing diverse outputs from incomplete or ambiguous inputs.
Given the data-intensive nature of deep learning, the construction and benchmarking of large-scale 3D shape datasets have become essential for training and evaluating generative models [CFG15, ZSK15].
However, learning to generate 3D shapes introduces unique challenges. Thus, the complexity and irregularity of 3D shapes and the different requirements of practical applications have resulted in the lack of a unified representation of 3D shapes [XLZ20]. Instead, various 3D representations are employed, each with its own structural properties, advantages, and limitations. These factors directly impact model design, data processing, and evaluation strategies.
We organize the discussion around three core components of the 3D shape generation pipeline: representations, generation methods, and evaluation protocols. Figure 3D Shape Generation: A Survey presents an overview of this survey. We first discuss the scope and related works in Section 2. Section 3 introduces major categories of 3D shape representations: explicit, implicit, and hybrid. Section 4 explores the most relevant 3D shape generation methods, with a focus on the feedforward generative models. Sections 5 and 6 present widely used datasets and metrics for assessing the fidelity and diversity of generated shapes, respectively. Finally, in Section 7, we discuss the open challenges and future directions.
@@ 2 Scope of this Survey
This survey aims to systematically summarize and categorize methods for 3D shape generation, together with the relevant representations, datasets, and metrics. The surveyed papers are primarily published in top-tier computer vision and computer graphics conferences and journals, as well as selected preprints published on arXiv in 2025. While it is challenging to exhaustively cover all methods related to 3D shape generation, we strive to include the most significant branches of the field. Rather than providing exhaustive explanations for each branch, we highlight representative works (as shown in Figure 1) to illustrate their main contributions and limitations. Further details can be found in the related work of these cited papers.
@@ 2.1 Related Survey
Our primary focus is on exploring techniques that generate 3D shapes. Therefore, this review does not encompass research on generation methods in the 2D domain. Several recent surveys have addressed specific aspects of 3D generation, including Deep Learning for 3D point cloud [GWH21], 3D-aware image synthesis [XX23], 3D generative models [SPX23], Text-to-3D [LZC24, LSC24], 3D generation methods based on scene representation [LZK24], implicit representations-based 3D shape generation [SWG24], and diffusion models for 3D generation [WPL25, Kum25]. In contrast, this survey provides an in-depth and comprehensive review of 3D shape generation methods and their related datasets and metrics (as shown in Table 1).
@@
3D shape representations are central to 3D computer vision and graphics [PJL21]. Shape representations can generally be categorized as either explicit or implicit. Explicit shape representations and learning algorithms depending on such representations directly parameterize the surface of the geometry, either as a point cloud [FSG17a, MYTB21, QSMG17, YHH19, WWG23, MWZ25], meshes [WZL18, JHK19, GC20], or surface patches [GFK18, WSS19].
While explicit representations are often computationally efficient and require few parameters, they tend to suffer from several limitations: discretization artifacts, difficulties in representing watertight surfaces, and constraints imposed by fixed or pre-defined topologies (meshes). In contrast, implicit representations define shapes as level sets of continuous functions over 3D space. These functions can be represented using discretized voxel grids [DRB18, WZX16] or, more recently, parameterized using neural networks, commonly referred to as neural implicit functions [YCWW22, BRS23].
@@ 3.1.1 Point Clouds
According to [PFS19], a point cloud is a lightweight 3D representation that closely matches the raw data that many sensors (e.g., LiDARs, depth cameras) provide, and hence is a natural fit for applying 3D learning.
@@
A primary limitation of point clouds is that they do not describe the topology and are not suitable for producing watertight surfaces [PFS19]. Another issue with point clouds as a representation is that they are unordered, which means that any permutation of a point set still describes the same shape. This complicates comparisons between two point sets, typically needed to define a loss function [ADMG18].
@@ 3.1.2 Meshes
According to [BKP10, BBL17, WZL18], a 3D mesh is a collection of vertices, faces, and edges that defines the shape of a 3D object; it can be represented by a graph structure (simplicial complex) $\mathcal{M}=(\mathcal{V},\mathcal{F},\mathcal{E})$, where
@@
$\mathcal{F}=\{f_{1},\dots,f_{F}\},\quad f_{i}\in\mathcal{V}\times\mathcal{V}\times\mathcal{V}$
is a set of triangular faces connecting them. However, it is more efficient to represent the connectivity of a triangle mesh in terms of the edges of the respective graph [BKP10],
$\mathcal{E}=\{e_{1},\dots,e_{E}\},\quad e_{i}\in\mathcal{V}\times\mathcal{V}.$
In this sense, a mesh-based representation stores the surface information cheaply as a list of vertices and faces that respectively define the geometric and topological information. Early approaches to mesh generation focused on predicting the parameters of category-specific mesh models [ZKB18, KPD19]. Although these methods produce manifold meshes, they are limited to object categories for which parameterized templates are available. Building upon this idea, works such as [GFK18, WZL18] extended mesh generation to a wider variety of object categories by incorporating topological priors, enabling more flexible and expressive 3D shape modeling beyond category-specific templates.
Compared to voxels, meshes are more compact and better suited to representing finer surface details. Compared to points, meshes are more controllable and exhibit better visual quality [GYW19].
@@ 3.1.3 Voxel Grids
The direct extension of pixels in 2D images to 3D is the voxel representation, which has a regular grid structure well-suited for convolutional neural networks (CNNs) [GFRG16, WZX16, GYW19].
@@
Moreover, as highlighted by [DRB18], small voxel sizes capture local detail but lack spatial context; large voxel sizes provide large spatial context but lack local detail.
@@ 3.2.1 Neural Radiance Fields (NeRF)
[MST22] introduce NeRF, a method for representing a 3D scene as an implicit function defined as
@@
where $\mathbf{x}$ is a 3D spatial coordinate, $\mathbf{d}$ is a 3D viewing direction, $\mathbf{c}$ is an RGB color, and $\sigma$ is a non-negative density value.
NeRFs model the appearance and geometry of 3D shapes by using volume rendering (ray matching) algorithms and positional encoding [MST22, MRP23]. This technique can efficiently reconstruct geometry given only a few multiview posed images. To achieve higher geometry quality, NeuS [WLL21] and VolSDF [YGKL21] extend radiance fields by using SDF instead of density as the geometry [SWG24].
@@ 3.2.2 Signed Distance Field (SDF)
As a function, SDF is more flexible than the common explicit representations, such as point clouds or meshes, and inherently allows topology manipulations such as constructive solid geometry (CSG) operations. Moreover, SDF allows the use of a technique known as “sphere tracing" [Har96], which can accelerate the rendering of path tracing [SWG24].
@@ 3.2.3 Neural Implicit Functions (NIFs)
Neural implicit functions (NIFs) have recently gained significant attention as a powerful and flexible framework for representing 3D scenes using neural networks. Unlike traditional explicit representations such as meshes, NIFs are not constrained by fixed spatial resolution, allowing them to capture fine geometric details at arbitrary scales [BRS23].
@@ 3.3.1 Point-Voxel
To address these limitations, [ILK21] proposes the use of point-voxel representations for 3D shape generation. This hybrid representation had been previously explored in methods such as PointCNN [LBS18], PV-RCNN [SGJ20], and PVD [ZDW21]. The core idea behind point-voxel is to voxelize the point clouds to enable efficient 3D convolutions while preserving point-level detail for finer geometric modeling.
@@ 3.3.2 Tri-Plane
In [CLC22], the Tri-Plane representation was introduced as a hybrid explicit-implicit 3D representation that balances computational efficiency with representational expressiveness.
In this formulation, features are explicitly arranged on three orthogonal, axis-aligned 2D planes–namely the $xy$, yzyz, and $xz$ planes, each with a resolution $N\times N\times C$, where $N$ denotes spatial resolution and $C$ the number of feature channels. As illustrated in Fig. 2, this structure contrasts with traditional 3D representations by encoding volumetric information through multiple planar projections.
Figure 2: Comparison of neural implicit, voxel-based, and tri-plane representations. (a) Neural implicit representations rely on fully connected (FC) layers with positional encoding (PE). They provide high fidelity but suffer from slow query times. (b) Explicit voxel grids, as well as hybrid variants with local decoders, allow for faster queries, yet their memory and computational costs grow rapidly with resolution. (c) Tri-plane representations strike a balance by supporting efficient queries while scaling more gracefully with resolution.
@@
In contrast to the original formulation, [SCP23] confirmed that simple addition (as shown in Equation 2) is an effective and efficient aggregation strategy for tri-plane features, avoiding the complexity of more elaborate fusion mechanisms.
The primary advantage of tri-planes representation lies in its efficiency–by keeping the decoder small and shifting the bulk of the expressive power into the explicit features. As demonstrated by [CLC22], this approach reduces the computational cost of neural rendering compared to fully implicit MLP architectures without losing expressiveness.
@@
According to [KYNS23], the first 3D generative models are based on GANs, learning a distribution of latents that can be decoded into various 3D representations such as point clouds [ADMG18, VFM18, SPK19] and implicit representations [KFW20, HASB20, CZ19, ILK21, ZLWT22]. Building upon this foundation, the rapid development of generative models introduces new generation methods such as flow-based models [YHH19, KBV20], which learn invertible mappings between complex and straightforward point distributions, enabling exact likelihood estimation and diverse point cloud synthesis; variational autoencoders (VAEs) [GFK18], which encode 3D shapes into a probabilistic latent space and reconstruct them via learned decoders, allowing interpolation and latent sampling; autoregressive models [SWL20, MGY19], which generate 3D geometry sequentially–either point-by-point or part-by-part–capturing structural dependencies; and, more recently, diffusion-based approaches [ZVW22, NJD22, Hea25], which progressively refine noisy inputs into clean 3D shapes through iterative denoising steps. Table 1 shows well-known examples of 3D generation methods using generative models. Table 1 also reports the metrics and datasets used to evaluate each generation method.
In this section, we explore the feedforward generation paradigm, proposed by [LZK24], which enables the direct synthesis of 3D representations through generative models.